# ShapeWords: Guiding Text-to-Image Synthesis with 3D Shape-Aware Prompts

### v1

@@ 1 Introduction
Recent advances in generative image models based on diffusion [17, 44, 37, 36, 2, 41, 38] have made it possible to generate impressive imagery from input text prompts. A challenge in text-to-image models has been to provide users with fine-grained control over shapes or forms in the synthesized images, which can be difficult to convey through text descriptions alone. To address this, conditioning methods have been proposed, such as ControlNet [51] and IP-adapter [48], that aim to capture the desired shape or form more explicitly through the use of edge or depth maps as input conditions.
@@
To overcome these challenges, we propose ShapeWords, a method designed to generate images that faithfully adhere to both the text prompt and a target 3D shape geometry, while at the same time allows users to easily explore both pose and shape variations. At the heart of ShapeWords is the embedding of the 3D shape into a shared space with textual descriptions, specifically OpenCLIP [10] space, enabling the target shape to be fully integrated with the prompt. Unlike view-specific conditioning (e.g., through depth, edge, or normal maps), our approach captures the entire 3D shape in the embedding, facilitating diverse image synthesis across challenging prompts, yet remaining geometrically consistent with the target 3D shape (Figure 1, top row). Furthermore, ShapeWords allows users to control the degree of deviation from the target shape – a feature that is particularly useful when the target structure is represented by coarse geometric primitives. This capability enables the generation of diverse, stylized images that retain plausible geometric variations of the specified target structure (Figure 1, bottom row).
@@ 2 Prior work
Denoising diffusion models [17, 44] have revolutionized image synthesis by generating high-quality, diverse, and plausible image content. To control the generation process, the most common condition is text. Popular diffusion approaches, such as Stable Diffusion [38], DALLE [37, 36, 2], Imagen [41] have used large-scale text-image datasets, including pretrained LLMs [6, 35, 11] for text-to-image synthesis. However, relying solely on text prompts cannot fully take advantage of the knowledge learned by diffusion models, especially when flexible and accurate control is needed in terms of form or layout. To this end, ControlNet [51] pioneered explicit conditioning through visual signals like depth or edge maps, enabling spatial control but limited to single viewpoints. T2I-Adapter [27] proposed a lighter-weight alternative through specialized adapters of visual control signals. UniControl [33] consolidated a wide array of controllable condition-to-image tasks within a single framework. IP-Adapter [48] proposed a decoupled cross-attention strategy for text features and image features from input conditions for more accurate controllable generation. In general, these methods rely on view-dependent control signals (e.g., depth, edge maps), lack explicit 3D geometry awareness, and may fail to generate images adhering to both the text prompt and target 3D shapes. In contrast, ShapeWords enables view-independent shape control, while adhering well both to target shape geometry and text prompts.
@@
Several approaches have explored novel view synthesis aimed at maintaining some level of 3D awareness for view consistency  [46, 23, 15, 24, 25]. However, these methods do not disentangle 3D structure from appearance and do not provide geometric control to users, meaning they generate images of a shape with consistent appearance across different views without user-driven control over geometry. A number of approaches have been proposed to incorporate structure guidance from depth maps or coarse 3D primitives disentangled from appearance. FreeControl [26] explores structure guidance in diffusion feature subspaces extracted from depth maps and other visual conditions. Ctrl-X [22] investigated more efficient, disentangled and zero-shot control of structure and appearance. LooseControl [4] introduced more flexible conditioning schemes through 3D scene boundary control, 3D box control and attribute editing. Diffusion Handles [29] enabled localized control of 3D object parts in diffusion models by introducing deformation handles to edit images while maintaining consistent perspective and structure across views. In our approach, we take a different route by embedding 3D shapes into tokens within text prompts, providing more explicit geometry guidance in image synthesis and combining it with additional specification of separate context, style, and appearance constraints through text.
@@
Several methods have explored embedding concepts into textual tokens for personalized image synthesis using textual inversion [13]. The required optimization for textual inversion is often very slow, thus various methods have explored more efficient fine-tuning strategies [19], feed-forward architectures to predict textual tokens [43, 21], or using hypernets [39]. Recently, more efficient methods of visual concept learning were introduced [40, 1, 14, 12] allowing for more efficient learning and transfer of visual concepts. A more spiritually similar approach to ours is that of “continuous 3D words”[9], which embeds 3D-aware attributes, such as time-of-day lighting, bird wing orientation, dolly zoom effects, and object pose, into learnable tokens. Viewpoint textual inversion[7] learns 3D view tokens that can be used to control the viewpoint for image synthesis. However, unlike these methods, our approach learns to embed 3D shapes directly into tokens, enabling image generation that is guided by both target 3D shape geometry and text. To our knowledge, text-to-image synthesis through learned 3D shape words has not been explored before.
@@ 3 Method
Given a text prompt ${\mathbf{p}}$ and a target 3D shape ${S}$, ShapeWords generates an image that reflects both the textual description and the desired shape. Users indicate the desired shape directly within the prompt using a special token, such as “a red [SHAPE-ID] on a beach”, where “SHAPE-ID” corresponds to a shape, e.g., one imported from a 3D shape database.
The pipeline (illustrated in Figure 2) proceeds as follows at test time. First, a shape representation is extracted from the chosen 3D shape using a pre-trained transformer, Point-BERT [49]. The text prompt is mapped into CLIP space through an OpenCLIP encoder [10], where the token “SHAPE-ID” is replaced by a category name for the shape (e.g., “chair”). Our method then applies a new module, named Shape2CLIP, trained to modify the prompt’s word embeddings, including the shape identifier token, so that the resulting prompt embedding integrates the desired 3D shape geometry in the prompt while preserving its original textual context. This shape-enhanced embedding is passed to a Stable Diffusion model [38], along with a user-defined parameter that controls the degree of 3D shape influence, to synthesize images consistent with both the textual and 3D shape cues.
In the following sections, we discuss preprocessing of the input shapes and text prompts to extract Point-Bert [49] and OpenCLIP [10] representations respectively (Section 3.1), then we discuss the Shape2CLIP module (Section 3.2), our inference pipeline (Section 3.3), and finally its training procedure (Section 3.4).
@@ 3.1 Preprocessing
To represent the input 3D shape, we use a pre-trained Point-BERT [49], a transformer-based model for point clouds. The input shape is first converted into a point cloud of $1024$ points using farthest point sampling, then passed through the Point-BERT architecture, which partitions the cloud into 64 patches. These patches are encoded by PointNet [32] and a transformer based encoder as a set of tokens representing the geometry of the shape. The resulting shape embedding, ${\mathbf{B}}\in\mathbb{R}^{65\times 384}$, consists of 65 tokens representing both patches and a class token in a 384-dimensional space. Trained via masked modeling for self-supervised learning, Point-BERT effectively captures structure-aware features demonstrated in various shape processing applications, such as part segmentation.
@@
As our base generative model, we use Stable Diffusion 2.1 whose text encoder is OpenCLIP ViT-H/14 [10]. Thus, we process the input textual prompt ${\mathbf{p}}$, through this OpenCLIP encoder to generate a sequence of $77$ token embeddings capturing the words in the prompt and their context in the prompt:
@@ 3.2 Shape2CLIP module
The residual function $\delta\mathbf{T}$ is constructed using cross-attention blocks [45]. In the first block, the shape representation ${\mathbf{B}}$ is treated as keys, while the original prompt embedding ${\mathbf{T}}$ serves as queries:
@@
followed by a MLP-based residual block [45]. In subsequent blocks, the updated prompt features act as queries, with Point-BERT representations serving as keys to further refine the prompt embedding. Our implementation uses a total of six cross-attention blocks. The output of last block is passed through a linear layer to yield the final modified embedding ${\mathbf{T}}^{\prime}$.
@@ 3.3 Guided Diffusion
The modified embedding ${\mathbf{T}}^{\prime}$ can be directly used as input to a diffusion model to generate images, formulated as ${\mathbf{I}}=D({\mathbf{z}},{\mathbf{T}}^{\prime})$, where ${\mathbf{z}}$ in the latent space and $D$ represents Stable Diffusion 2.1 [38] in our implementation. Notably, our method does not require any additional shape conditions (e.g., depth, normal, or edge maps).
@@
As shown in our results, varying $\lambda$ from 00 to $1$ gradually shifts from disregarding the shape influence to incorporating it in the generated images. Intermediate values yield plausible images, while higher $\lambda$ values produce shapes in the generated images that increasingly match the desired target shape.
@@ 3.4 Training
To train the Shape2CLIP module, we constructed a dataset of shape-prompt-image triplets based on ShapeNet [8] reference shapes. Images were generated using ControlNet [51] conditioned on ShapeNet’s depth maps, as provided by the ULIP authors [47], where each shape is rendered from 30 viewpoints, rotated in 12-degree increments around the vertical axis with fixed elevation (see [47] for details).
For each depth image, we applied a randomly selected prompt from a set of $13716$ prompts for additional ControlNet conditioning, created from a base set of $100$ prompts, then augmented with variants produced by ChatGPT [28]. To achieve diversity and structural agnosticism in prompts, this set was created by combining $127$ artistic mediums (e.g., “painting,” “watercolor,” “sketch”) with $108$ style adjectives (e.g., “colorful,” “pixelated,” “fantasy”), deliberately avoiding references to specific 3D structures to help the model learn generalizable mappings instead of overfitting to particular geometries or appearance combinations. To reduce bias and enhance background diversity, we used the Stable Diffusion’s inpainting model [38] to modify the backgrounds while preserving the foreground objects. The above procedure resulted in a diverse dataset of $1.58$M prompt-image pairs (30 per each ShapeNet shape), generated from all depth images in the training split from 3DILG [50] in ShapeNet. As demonstrated in Section 4, although our method is trained on data generated by ControlNet, ShapeWords exhibits strong generalization capabilities. Notably, ShapeWords achieves significantly better performance on compositional prompts, a setting that ControlNet struggles to handle effectively. The dataset will be publicly released along with our source code upon acceptance.
@@
We train our Shape2CLIP model using the Score Distillation Sampling (SDS) loss [31]. Specifically, for sampled noise ${\mathbf{\epsilon}}_{t,i}\sim\mathcal{N}(0,\,\mathbf{I})$ for a training image $i$ at step $t$, we use the pretrained Stable Diffusion model to predict the noise $\hat{{\mathbf{\epsilon}}}_{t,i}$, and backpropagate the loss:
@@
where $W(t)$ is a time-dependent weighting function proposed in DreamTime [18] for enhancing the stability of the SDS-based training. We provide details in supplementary.
@@ 4 Evaluation
(a) Prompt Adherence: We evaluate how well the images generated by each competing method align with the given prompt. We use the standard score of CLIP similarity [34].
(b) Shape Adherence: We assess how well the generated images from each method adhere to a reference 3D shape. To this end, we compare the shapes in the generated images with the reference 3D shapes based on their silhouette. To extract silhouttes from reference 3D shapes, we render them from a target pose, extract the silhouette, then we compare it with the silhouette of the shape in the generated images conditioned on the target pose. The silhouette from generated images are extracted through the foreground detection model [52]. The similarity between silhouettes is measured using standard geometric similarity metrics, specifically Intersection over Union (S-IOU) and Chamfer Distance (S-CD), averaged per shape across six uniform views sampled with 60 degree increments, starting at 0.
(c) Image Plausibility: We evaluate the aesthetic quality of the generated images using the aesthetics score (Aes.) [42]. We also include the commonly used image quality generation metrics of FID [16] and KID [5]. Given reported issues with Inception-based features [20, 30], we compute FID and KID on CLIP features as recommended in [3].
@@
(a) ControlNet: ControlNet receives two inputs: (i) a depth image rendered from a target 3D shape in a specified pose (we use inverted depth images provided by the ULIP project [47]), and (ii) a text prompt from our test datasets where [SHAPE-ID] is replaced with the ShapeNet category label at the finest hierarchical level of ShapeNet. ControlNet’s control strength is set to 1 (default). This setup makes ControlNet a strong baseline for shape adherence as it generates images whose object silhouettes closely match those of the input depth image. However, with compositional prompts, we observed that ControlNet tends to prioritize depth image content over textual context, leading to weaker prompt adherence.
@@
Our qualitative evaluation (Figure 5) also suggests that ControlNet’s depth conditioning tends to prioritize shape adherence at the expense of prompt adherence and aesthetic plausibility in the generated images especially for complex prompts. We also include a qualitative comparison with the recent method of Ctrl-X [22] in the same figure, where we observed similar, or even worse behavior in terms of shape adherence compared to ControlNet. In contrast, our method strikes a much better balance, achieving a more effective trade-off between maintaining shape fidelity and adhering to the input prompt while producing visually plausible results.
@@
Figure 8: Ablation on guidance strength $\lambda$. Here we vary $\lambda$ from 00 to $1$ for the prompt “aquarelle drawing of a [SHAPE-ID]” (random seed is fixed, so results for $\lambda=0$ are the same). Increasing level of guidance strength increases resemblance to target shape. The intermediate results still remain aesthetically plausible.

### v2

@@ 1 Introduction
Recent advances in generative image models based on diffusion [ddpm, ddim, dalle1, dalle2, dalle3, imagen, rombach2022stable] have made it possible to generate impressive imagery from input text prompts. A challenge in text-to-image models has been to provide users with fine-grained control over shapes or forms in the synthesized images, which can be difficult to convey through text descriptions alone. To address this, conditioning methods have been proposed, such as ControlNet [zhang2023controlnet] and IP-adapter [ye2023ip-adapter], that aim to capture the desired shape or form more explicitly through the use of edge or depth maps as input conditions.
@@
To overcome these challenges, we propose ShapeWords, a method designed to generate images that faithfully adhere to both the text prompt and a target 3D shape geometry, while at the same time allows users to easily explore both pose and shape variations. At the heart of ShapeWords is the embedding of the 3D shape into a shared space with textual descriptions, specifically OpenCLIP [cherti2023OpenCLIP] space, enabling the target shape to be fully integrated with the prompt. Unlike view-specific conditioning (e.g., through depth, edge, or normal maps), our approach captures the entire 3D shape in the embedding, facilitating diverse image synthesis across challenging prompts, yet remaining geometrically consistent with the target 3D shape (Figure 1, top row). Furthermore, ShapeWords allows users to control the degree of deviation from the target shape – a feature that is particularly useful when the target structure is represented by coarse geometric primitives. This capability enables the generation of diverse, stylized images that retain plausible geometric variations of the specified target structure (Figure 1, bottom row).
@@ 2 Prior work
Denoising diffusion models [ddpm, ddim] have revolutionized image synthesis by generating high-quality, diverse, and plausible image content. To control the generation process, the most common condition is text. Popular diffusion approaches, such as Stable Diffusion [rombach2022stable], DALLE [dalle1, dalle2, dalle3], Imagen [imagen] have used large-scale text-image datasets, including pretrained LLMs [gpt3, t5, bert] for text-to-image synthesis. However, relying solely on text prompts cannot fully take advantage of the knowledge learned by diffusion models, especially when flexible and accurate control is needed in terms of form or layout. To this end, ControlNet [zhang2023controlnet] pioneered explicit conditioning through visual signals like depth or edge maps, enabling spatial control but limited to single viewpoints. T2I-Adapter [mou2023t2i] proposed a lighter-weight alternative through specialized adapters of visual control signals. UniControl [qin2023unicontrol] consolidated a wide array of controllable condition-to-image tasks within a single framework. IP-Adapter [ye2023ip-adapter] proposed a decoupled cross-attention strategy for text features and image features from input conditions for more accurate controllable generation. In general, these methods rely on view-dependent control signals (e.g., depth, edge maps), lack explicit 3D geometry awareness, and may fail to generate images adhering to both the text prompt and target 3D shapes. In contrast, ShapeWords enables view-independent shape control, while adhering well both to target shape geometry and text prompts.
@@
Several approaches have explored novel view synthesis aimed at maintaining some level of 3D awareness for view consistency  [xiang20233daware, liu2023zero123, gu2022stylenerf, liu2023syncdreamer, melaskyriazi2023realfusion]. However, these methods do not disentangle 3D structure from appearance and do not provide geometric control to users, meaning they generate images of a shape with consistent appearance across different views without user-driven control over geometry. A number of approaches have been proposed to incorporate structure guidance from depth maps or coarse 3D primitives disentangled from appearance. FreeControl [mo024freecontrol] explores structure guidance in diffusion feature subspaces extracted from depth maps and other visual conditions. Ctrl-X [lin2024ctrlx] investigated more efficient, disentangled and zero-shot control of structure and appearance. LooseControl [bhat2024loosecontrol] introduced more flexible conditioning schemes through 3D scene boundary control, 3D box control and attribute editing. Diffusion Handles [pandey2024diffusionhandles] enabled localized control of 3D object parts in diffusion models by introducing deformation handles to edit images while maintaining consistent perspective and structure across views. In our approach, we take a different route by embedding 3D shapes into tokens within text prompts, providing more explicit geometry guidance in image synthesis and combining it with additional specification of separate context, style, and appearance constraints through text.
@@
Several methods have explored embedding concepts into textual tokens for personalized image synthesis using textual inversion [gal2023textual]. The required optimization for textual inversion is often very slow, thus various methods have explored more efficient fine-tuning strategies [kumari2023custom], feed-forward architectures to predict textual tokens [shi2024instantbooth, li2024photomaker], or using hypernets [ruiz2024hyperdreambooth]. Recently, more efficient methods of visual concept learning were introduced [safaee2024clic, avrahami2023break, gal2023encoder, ding2024clip] allowing for more efficient learning and transfer of visual concepts. A more spiritually similar approach to ours is that of “continuous 3D words”[cheng2024learning], which embeds 3D-aware attributes, such as time-of-day lighting, bird wing orientation, dolly zoom effects, and object pose, into learnable tokens. Viewpoint textual inversion[viewneti] learns 3D view tokens that can be used to control the viewpoint for image synthesis. However, unlike these methods, our approach learns to embed 3D shapes directly into tokens, enabling image generation that is guided by both target 3D shape geometry and text. To our knowledge, text-to-image synthesis through learned 3D shape words has not been explored before.
@@ 3 Method
Given a text prompt $\mathbf{p}$} and a target 3D shape S{S}, ShapeWords generates an image that reflects both the textual description and the desired shape. Users indicate the desired shape directly within the prompt using a special token, such as “a red [SHAPE-ID] on a beach”, where “SHAPE-ID” corresponds to a shape, e.g., one imported from a 3D shape database.
The pipeline (illustrated in Figure 2) proceeds as follows at test time. First, a shape representation is extracted from the chosen 3D shape using a pre-trained transformer, Point-BERT [yu2022pointbert]. The text prompt is mapped into CLIP space through an OpenCLIP encoder [cherti2023OpenCLIP], where the token “SHAPE-ID” is replaced by a category name for the shape (e.g., “chair”). Our method then applies a new module, named Shape2CLIP, trained to modify the prompt’s word embeddings, including the shape identifier token, so that the resulting prompt embedding integrates the desired 3D shape geometry in the prompt while preserving its original textual context. This shape-enhanced embedding is passed to a Stable Diffusion model [rombach2022stable], along with a user-defined parameter that controls the degree of 3D shape influence, to synthesize images consistent with both the textual and 3D shape cues.
In the following sections, we discuss preprocessing of the input shapes and text prompts to extract Point-Bert [yu2022pointbert] and OpenCLIP [cherti2023OpenCLIP] representations respectively (Section 3.1), then we discuss the Shape2CLIP module (Section 3.2), our inference pipeline (Section 3.3), and finally its training procedure (Section 3.4).
@@ 3.1 Preprocessing
To represent the input 3D shape, we use a pre-trained Point-BERT [yu2022pointbert], a transformer-based model for point clouds. The input shape is first converted into a point cloud of $1024$ points using farthest point sampling, then passed through the Point-BERT architecture, which partitions the cloud into 64 patches. These patches are encoded by PointNet [qi2017pointnet] and a transformer based encoder as a set of tokens representing the geometry of the shape. The resulting shape embedding, ${\mathbf{B}}\in\mathbb{R}^{65\times 384}$, consists of 65 tokens representing both patches and a class token in a 384-dimensional space. Trained via masked modeling for self-supervised learning, Point-BERT effectively captures structure-aware features demonstrated in various shape processing applications, such as part segmentation.
@@
As our base generative model, we use Stable Diffusion 2.1 whose text encoder is OpenCLIP ViT-H/14 [cherti2023OpenCLIP]. Thus, we process the input textual prompt $\mathbf{p}$}, through this OpenCLIP encoder to generate a sequence of $77$ token embeddings capturing the words in the prompt and their context in the prompt:
@@ 3.2 Shape2CLIP module
The residual function $\delta\mathbf{T}$ is constructed using cross-attention blocks [vaswani2017attention]. In the first block, the shape representation $\mathbf{B}$} is treated as keys, while the original prompt embedding $\mathbf{T}$} serves as queries:
@@
followed by a MLP-based residual block [vaswani2017attention]. In subsequent blocks, the updated prompt features act as queries, with Point-BERT representations serving as keys to further refine the prompt embedding. Our implementation uses a total of six cross-attention blocks. The output of last block is passed through a linear layer to yield the final modified embedding $\mathbf{T}$$^{\prime}$.
@@ 3.3 Guided Diffusion
The modified embedding $\mathbf{T}$$^{\prime}$ can be directly used as input to a diffusion model to generate images, formulated as ${\mathbf{I}}=D({\mathbf{z}},{\mathbf{T}}^{\prime})$, where $\mathbf{z}$} in the latent space and $D$ represents Stable Diffusion 2.1 [rombach2022stable] in our implementation. Notably, our method does not require any additional shape conditions (e.g., depth, normal, or edge maps).
@@
As shown in our results, varying $\lambda$ from 0 to $1$ gradually shifts from disregarding the shape influence to incorporating it in the generated images. Intermediate values yield plausible images, while higher $\lambda$ values produce shapes in the generated images that increasingly match the desired target shape.
@@ 3.4 Training
To train the Shape2CLIP module, we constructed a dataset of shape-prompt-image triplets based on ShapeNet [chang2015shapenet] reference shapes. Images were generated using ControlNet [zhang2023controlnet] conditioned on ShapeNet’s depth maps, as provided by the ULIP authors [xue2023ulip], where each shape is rendered from 30 viewpoints, rotated in 12-degree increments around the vertical axis with fixed elevation (see [xue2023ulip] for details).
For each depth image, we applied a randomly selected prompt from a set of $13716$ prompts for additional ControlNet conditioning, created from a base set of $100$ prompts, then augmented with variants produced by ChatGPT [chatgpt]. To achieve diversity and structural agnosticism in prompts, this set was created by combining $127$ artistic mediums (e.g., “painting,” “watercolor,” “sketch”) with $108$ style adjectives (e.g., “colorful,” “pixelated,” “fantasy”), deliberately avoiding references to specific 3D structures to help the model learn generalizable mappings instead of overfitting to particular geometries or appearance combinations. To reduce bias and enhance background diversity, we used the Stable Diffusion’s inpainting model [rombach2022stable] to modify the backgrounds while preserving the foreground objects. The above procedure resulted in a diverse dataset of 1.581.58M prompt-image pairs (30 per each ShapeNet shape), generated from all depth images in the training split from 3DILG [zhang2022dilg] in ShapeNet. As demonstrated in Section 4, although our method is trained on data generated by ControlNet, ShapeWords exhibits strong generalization capabilities. Notably, ShapeWords achieves significantly better performance on compositional prompts, a setting that ControlNet struggles to handle effectively. The dataset will be publicly released along with our source code upon acceptance.
@@
We train our Shape2CLIP model using the Score Distillation Sampling (SDS) loss [poole2023dreamfusion]. Specifically, for sampled noise ${\mathbf{\epsilon}}_{t,i}\sim\mathcal{N}(0,\,\mathbf{I})$ for a training image $i$ at step $t$, we use the pretrained Stable Diffusion model to predict the noise $\hat{{\mathbf{\epsilon}}}_{t,i}$, and backpropagate the loss:
@@
where $W(t)$ is a time-dependent weighting function proposed in DreamTime [huang2024dreamtime] for enhancing the stability of the SDS-based training. We provide details in supplementary.
@@ 4 Evaluation
(a) Prompt Adherence: We evaluate how well the images generated by each competing method align with the given prompt. We use the standard score of CLIP similarity [radford2021learning].
(b) Shape Adherence: We assess how well the generated images from each method adhere to a reference 3D shape. To this end, we compare the shapes in the generated images with the reference 3D shapes based on their silhouette. To extract silhouttes from reference 3D shapes, we render them from a target pose, extract the silhouette, then we compare it with the silhouette of the shape in the generated images conditioned on the target pose. The silhouette from generated images are extracted through the foreground detection model [zheng2024birefnet]. The similarity between silhouettes is measured using standard geometric similarity metrics, specifically Intersection over Union (S-IOU) and Chamfer Distance (S-CD), averaged per shape across six uniform views sampled with 60 degree increments, starting at 0.
(c) Image Plausibility: We evaluate the aesthetic quality of the generated images using the aesthetics score (Aes.) [Schuhmann]. We also include the commonly used image quality generation metrics of FID [heusel2017gans] and KID [binkowski2018demystifying]. Given reported issues with Inception-based features [kynkaanniemi2022role, parmar2022aliased], we compute FID and KID on CLIP features as recommended in [betzalel2022study].
@@
(a) ControlNet: ControlNet receives two inputs: (i) a depth image rendered from a target 3D shape in a specified pose (we use inverted depth images provided by the ULIP project [xue2023ulip]), and (ii) a text prompt from our test datasets where [SHAPE-ID] is replaced with the ShapeNet category label at the finest hierarchical level of ShapeNet. ControlNet’s control strength is set to 1 (default). This setup makes ControlNet a strong baseline for shape adherence as it generates images whose object silhouettes closely match those of the input depth image. However, with compositional prompts, we observed that ControlNet tends to prioritize depth image content over textual context, leading to weaker prompt adherence.
@@
Our qualitative evaluation (Figure 5) also suggests that ControlNet’s depth conditioning tends to prioritize shape adherence at the expense of prompt adherence and aesthetic plausibility in the generated images especially for complex prompts. We also include a qualitative comparison with the recent method of Ctrl-X [lin2024ctrlx] in the same figure, where we observed similar, or even worse behavior in terms of shape adherence compared to ControlNet. In contrast, our method strikes a much better balance, achieving a more effective trade-off between maintaining shape fidelity and adhering to the input prompt while producing visually plausible results.
@@
Figure 8: Ablation on guidance strength $\lambda$. Here we vary $\lambda$ from 0 to 11 for the prompt “aquarelle drawing of a [SHAPE-ID]” (random seed is fixed, so results for $\lambda=0$ are the same). Increasing level of guidance strength increases resemblance to target shape. The intermediate results still remain aesthetically plausible.